Erin starts riding her bike. She is travelling at 10km/h. An hour after Erin starts, her sister, Makayla starts riding her bike. She is travelling at 16km/h. After how long (time) will Makayla catch up to her sister?
Answer by mananth(16946)

Erin 10 km/h
Makyala 16 km/h
Difference in time= 1.00 hour
Erin will have covered 10.00 km when Makyala starts
catch up distance= 10.00 km
catch up speed = 16 -10
catch up speed = 6 km/h
Catchup time = catchup distance/catch up speed
catch up time= 10/6 = 5/3 hours